OpenStudy (kmullis6):

PLEASE HELP ON ALGEBRA 1 I DONT GET THIS STUFF /: Will fan and give medal! :D

OpenStudy (kmullis6):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (anonymous):

They have the same number of solutions. They are both linear, absolute value problems, so they each have 2 solution. To solve you have to solve for 2 equations. The original equation, and then the negative of that equation. The reason for this is because if we absolute value the negative of something, it becomes positive, leading to 2 possible solutions. For example: if we have absolute value of x = 1 The solutions are 1 and -1. Moving on. Let's look at 1 problem at a time. Set up the first problem: 5x + 6 = 41 => 5x = 35 => x = 7 -5x - 6 = 41 => -5x = 47 => x = -47/5 Do the same for the second problem: 2x + 13 = 28 => 2x = 15 => x = 15/2 -2x -13 = 28 => -2x = 41 => x = -41/2

OpenStudy (kmullis6):

@IAmSinged so it's A?

OpenStudy (anonymous):

B. They have the same # of solutions.

OpenStudy (kmullis6):

Ohhh ok thanks so much for your help i really appreciate it (: